# Hierarchical Object-Oriented POMDP Planning for Object Rearrangement

### v2

@@ 7 Conclusion and future work
In this paper, we presented a novel Hierarchical Object-Oriented POMDP Planner (HOO-POMDP) for solving multi-object rearrangement problems in partially observable, multi-room environments. Our approach decomposes the complex task into a high-level abstract POMDP planner for generating sub-goals and low-level policies for execution. Key components include an object-oriented state representation, belief updating to handle perception uncertainty, and an abstraction system to bridge the gap between continuous and discrete planning. Experimental results across multiple datasets demonstrate the effectiveness of our approach in handling challenging scenarios such as blocked paths and goals. The HOO-POMDP framework showed robust performance in terms of success rate and efficiency comparable to Oracle baselines with perfect knowledge or perfect detection. Notably, our method scaled well to environments with more objects and rooms. One of the ways to expand the scope is to relax the assumption of object independence partially. We can allow objects to be dependent on a small number of objects (e.g., objects in their close vicinity). Belief updates can now consider a small set of objects at any time. This relaxation helps maintain the efficient belief update while accounting for more real-world situations such as object-object interaction. Another potential future work is to handle stacking of objects and more cramped spaces, where more careful reasoning about object interactions is needed to plan the actions and order them appropriately.
@@ A.1.1 Detection Model : YoloV10
We collect data from the AI2Thor simulator. We do this by placing the agent in random locations in 500 scenes and extracting the RGB images along with the ground truth object bounding box annotations from the simulator. We have $50$ pickupable object classes in all our scenes combined. We train the YoloV10 detector on 10,000 images collected from these 500 scenes.
@@ A.1.2 Observation Probability for POMDP
$\Pr(z_{i}|s_{target_{i}},s_{\text{r}})=\begin{cases}1.0-\text{TP}&s_{i}\in\mathcal{V}(s_{\text{robot}})\wedge z_{i}=\text{null}\\ \delta\text{FP}/|\mathcal{V}_{E}(r)|&s_{i}\in\mathcal{V}(s_{\text{robot}})\wedge\|z_{i}-s_{i}\|>3\sigma\\ \delta z_{i}&s_{i}\in\mathcal{V}(s_{\text{robot}})\wedge\|z_{i}-s_{i}\|\leq 3\sigma\\ 1.0-\text{FP}&s_{i}\notin\mathcal{V}(s_{\text{robot}})\wedge z_{i}=\text{null}\\ \delta\text{FP}/|\mathcal{V}_{E}(r)|&s_{i}\notin\mathcal{V}(s_{\text{robot}})\wedge z_{i}\neq\text{null}\end{cases}$
@@
The list of $TP$, $FP$ and $r$ for the object classes in the dataset presented in table 2.
@@ A.2.1 Overview
The dataset consists of 400 distinct room configurations, with varying complexity in terms of room count and object arrangements.
A.2.2 Dataset Composition
@@
  • Object Types: Comprehensive selection from AI2Thor environment (see Appendix).
@@
A.2.3 Room Configuration Distribution
@@
A.2.4 Object Placement Criteria
@@ A.3 POUCT
1 Function POUCTPlanner(abstractState, beliefState):
2       $b\leftarrow beliefState$;
3       $abs\leftarrow abstractState$;
4       $T\leftarrow\{\}$;
5       for $j=0$ to $SIMULATIONS$ do
6             $\hat{s}\leftarrow$ SAMPLE($b$);
7             SIMULATE($\hat{s},\{\},0$,abs);
8
9      $abstractAction\leftarrow\arg\max_{a}V(ha)$;
10       return $abstractAction$;
11
12Function SAMPLE($b$):
13       for $o\in Obj$ do
14             $\hat{s}_{o}\sim b_{o}$;
15
16      return $\bigcup\hat{s}_{o}$;
17
18Function SIMULATE($s,h,depth,abs$):
19       if $\gamma^{depth}<\epsilon$ then
20             return 0;
21
22      if $h\notin T$ then
23             for all $a\in\mathcal{A}$ do
24                   $T(ha)\leftarrow\langle N_{init}(ha),V_{init}(ha)\rangle$;
25
26            return ROLLOUT$(s,h,depth,abs)$;
27
28      $a\leftarrow$ selectMaxAction();
29       $(s^{\prime},z,r)\sim\mathcal{G}(s,a)$;
30       $R\leftarrow r+\gamma\cdot$ SIMULATE($s^{\prime},hao,depth+1$);
31       $T(ha)\leftarrow\left\langle N(ha)+1,V(ha)+\frac{R-V(ha)}{N(ha)}\right\rangle$;
32       return $R$;
33
34Function ROLLOUT($s,h,depth,abs$):
35       if $\gamma^{depth}<\epsilon$ then
36             return 0;
37
38      $a\sim\pi_{rollout}(h,abs)$;
39       $(s^{\prime},o,r)\sim\mathcal{G}(s,a)$;
40       return $r+\gamma\cdot$ ROLLOUT($s^{\prime},hao,depth+1$);
41
Algorithm 3 POUCT Planner (Modified version from Silver & Veness (2010))
@@ A.4 Qualitative Results
We test our system with different depths for the MCTS planner to see how much look-ahead affects the performance of our system. MCTS search depth is one of the important factors determining the amount of exploration and the time taken for search at step. Results are shown in table 3. The depth for the system - OURS is $12$, and the depth for OURS-MCTS_1 is $1$. From the table, we can see that the greedy approach of just looking ahead by 1 step is not enough to solve the rearrangement problem. This is because, when we look ahead only 1 step, we do not get enough reward/feedback from the world about what is a good path to take and hence makes it extremely hard to solve the problem. We can also see that, in the problems it does solve, it takes significantly longer to solve. We did not compute the results for 15 and 20 objects in this setting as it failed to solve any problems in 3-4 rooms with ten objects.
Dataset Objs #BG #Sw #BP #Rm #V Ours OURS-MCSTS_1
@@
2 1 1 2 2-3 31 53 398 4.52 1 10 1102
Multi RoomR 10 1 1 0 2 2-3 32 65 710 7.89 0 7 0
2 1 1 2 2-3 21 49 789 8.98 0 3 0
10 2 1 0 3-4 1-2 30 62 1189 11.45 0 0 0
2 1 1 3-4 1-2 18 44 1321 12.97 0 0 0
15 1 1 0 3-4 2-3 22 59 1228 16.89 NC NC NC
2 1 1 3-4 2-3 14 41 1416 17.12 NC NC NC
20 2 1 0 3-4 2-4 17 55 1621 21.61 NC NC NC
2 1 1 3-4 2-4 10 36 1786 23.79 NC NC NC
Table 3: Performance metrics for our method across different depths for MCTS search. Metrics: Success Score (SS), Object Success Rate (OSR), Task Actions (TA), execution Time in minutes, and number of objects initially visible (#V). The difficulty parameters include the number of blocked goals (#BG), objects to be swapped (#Sw), blocking objects (#BP), and number of rooms (#Rm). NC: Not computed.

### v3

@@ 7 Conclusion and future work
In this paper, we presented a novel Hierarchical Object-Oriented POMDP Planner (HOO-POMDP) for solving multi-object rearrangement problems in partially observable, multi-room environments. Our approach decomposes the complex task into a high-level abstract POMDP planner for generating sub-goals and low-level policies for execution. Key components include an object-oriented state representation, belief update that handles perception uncertainty, and an abstraction system that bridges the gap between continuous and discrete planning. Experimental results across multiple datasets demonstrate the effectiveness of our approach in handling challenging scenarios such as blocked paths and goals. The HOO-POMDP framework showed robust performance in terms of success rate and efficiency comparable to oracle baselines with perfect knowledge or perfect detection. Notably, our method scaled well to environments with more objects and rooms.
Future work could partially relax our object independence assumption to better reflect real-world scenarios and local object dependencies. Another key direction for future work is deploying our framework on physical robots by replacing the reinforcement learning controller with a low-level robot controller.
@@ A.1.1 Detection Model : YoloV10
We collect data from the AI2Thor simulator. We do this by placing the agent in random locations in 500 scenes and extracting the RGB images along with the ground truth object bounding box annotations from the simulator. We have $50$ pickupable object classes in all our scenes combined. We train the YoloV10 detector (YoloV10-Large, with 25 million parameters) on 10,000 images collected from these 500 scenes. It is trained for 500 epochs, batch size 16. It is trained on RTX 3080 for 12 hours.
@@ A.1.2 Observation Probability for POMDP
$\Pr(z_{i}|s_{target_{i}},s_{\text{robot}})=\begin{cases}1.0-\text{TP}&s_{target_{i}}\in\mathcal{V}(s_{\text{robot}})\wedge z_{i}=\text{null}\\ \delta\text{FP}/|\mathcal{V}_{E}(r)|&s_{target_{i}}\in\mathcal{V}(s_{\text{robot}})\wedge\|z_{i}-s_{target_{i}}\|>3\sigma\\ \delta*TP&s_{target_{i}}\in\mathcal{V}(s_{\text{robot}})\wedge\|z_{i}-s_{target_{i}}\|\leq 3\sigma\\ 1.0-\text{FP}&s_{target_{i}}\notin\mathcal{V}(s_{\text{robot}})\wedge z_{i}=\text{null}\\ \delta\text{FP}/|\mathcal{V}_{E}(r)|&s_{target_{i}}\notin\mathcal{V}(s_{\text{robot}})\wedge z_{i}\neq\text{null}\end{cases}$
@@
The list of $TP$, FPFP, and $r$ for the object classes in the dataset presented in table 2.
@@ A.2.1 Overview
Recent advancements in robotic rearrangement have been facilitated by datasets such as RoomR Batra et al. [2020] and ProcThorRearrangement. However, these datasets exhibit significant limitations that prevent them from capturing the complexity of real-world rearrangement scenarios. Specifically, existing datasets are constrained in their scope, typically featuring only 5 objects and limited room configurations (single room in RoomR, two rooms in ProcThorRearrangement). These constraints fail to represent the challenges inherent in real-world environments, where rearrangement tasks frequently span multiple rooms and involve numerous objects.
To address these limitations, we present MultiRoomR, a comprehensive dataset designed to bridge three critical gaps in existing benchmarks. First, MultiRoomR emphasizes partial observability by incorporating scenes with 2-4 rooms where most objects are not immediately visible, thereby ensuring that systems must develop robust strategies for handling incomplete information. Second, the dataset increases scene complexity by including 10-20 objects per environment, necessitating the development of efficient and scalable solutions that can handle larger object sets without compromising optimality. Third, MultiRoomR introduces realistic constraints such as blocked paths, with 50% of scenes containing at least one object that obstructs access to other objects. This feature specifically tests a system’s ability to reason about sequential manipulation, as encountered in practical scenarios where intermediate object movements are necessary to complete the primary rearrangement task. By incorporating these challenging aspects, MultiRoomR provides a more rigorous evaluation framework that better aligns with real-world requirements. Systems that demonstrate strong performance on this dataset are likely to be more suitable for deployment in actual home environments, where partial observability, multi-object manipulation, and complex spatial reasoning are commonplace challenges. We also provide the dataset  [^1] Dataset submitted as part of supplementary material and code  [^2] Link to code: https://anonymous.4open.science/r/ICML_POMDP_rearrangement-4460/ for the community to use. The dataset consists of 400 distinct room configurations, with varying complexity in terms of room count and object arrangements. The dataset composition and room configurations are presented in the rest of this section.
A.2.2 Dataset Creation Criteria
Data generation has 2 parts:
  • Basic scene generation is done using ProcThor [Deitke et al., 2022]. To increase task complexity, we (1) add the extra condition, the average distance between an object and its goal is above 25 steps, and (2) sample goal positions in different rooms for at least 50% of the objects.
  • Second, for generating blocked path scenes, we use the connected components algorithm. We construct a connected graph of all 2D navigable points in the room. We then find 2*2 grid locations that if not navigable will make a portion of the house not reachable. Among these possible options, we pick the grid that blocks the maximum number of objects from the current starting location of the agent to ensure that not moving a blocking object (of size 2*2) renders a large number of objects inaccessible. We will add this in the appendix.
A.2.3 Dataset Composition
Figure 3: Range of percentage of objects visible at the beginning of a scene
@@ A.2.3 Dataset Composition
  • Object Types: Comprehensive selection from AI2Thor environment (see Table 2)
@@
  • Figure 3 shows the range of percentage of objects visible at the beginning of the scene. Lower the objects visible, more exploration needs to be done thus making the problem harder. It ranges between 10-40% in our dataset where it is 60% in Mirakhor et al. [2024b] Table 1. This implies, only 20-30 % of objects are visible in the beginning for two-room scenes and so on for other room settings.
A.2.4 Room Configuration Distribution
@@
  • Initial visibility : 20-30% of objects visible at the beginning of the scene
@@
    • Initial visibility : 10-20% of objects visible at the beginning of the scene
@@
    • Initial visibility : 10-20% of objects visible at the beginning of the scene
A.2.5 Object Placement Criteria
@@ A.3 POUCT
  1. 1.
    Input: abstractState absabs, beliefState $b$
  2. 2.
    $T\leftarrow\{\}$
  3. 3.
    For $j=0$ to $SIMULATIONS$:
  4. 4.
    $\hat{s}\leftarrow$ SAMPLE($b$)
  5. 5.
    SIMULATE($\hat{s},\{\},0,abs$)
  6. 6.
    Return $\arg\max_{a}V(ha)$
  7. 7.
    Function SAMPLE($b$):
  8. 8.
    For each object $o$:
  9. 9.
    Sample $\hat{s}_{o}\sim b_{o}$
  10. 10.
    Return $\bigcup\hat{s}_{o}$
  11. 11.
    Function SIMULATE($s,h,depth,abs$):
  12. 12.
    if $\gamma^{depth}<\epsilon$: return 0
  13. 13.
    if $h\notin T$:
  14. 14.
    Initialize $T(ha)$ for all actions $a$
  15. 15.
    Return ROLLOUT($s,h,depth,abs$)
  16. 16.
    $a\leftarrow$ selectMaxAction()
  17. 17.
    $(s^{\prime},z,r)\sim\mathcal{G}(s,a)$
  18. 18.
    $R\leftarrow r+\gamma\cdot$ SIMULATE($s^{\prime},hao,depth+1$)
  19. 19.
    Update $T(ha)$ with new value and count
  20. 20.
    Return $R$
  21. 21.
    Function ROLLOUT($s,h,depth,abs$):
  22. 22.
    if $\gamma^{depth}<\epsilon$: return 0
  23. 23.
    $a\sim\pi_{rollout}(h,abs)$
  24. 24.
    $(s^{\prime},o,r)\sim\mathcal{G}(s,a)$
  25. 25.
    Return $r+\gamma\cdot$ ROLLOUT($s^{\prime},hao,depth+1$)
Algorithm 3 POUCT Planner
@@ A.4 Qualitative Results
We test our system with different depths for the MCTS planner to see how much look-ahead affects the performance of our system. MCTS search depth is one of the important factors determining the amount of exploration and the time taken for search at step. Results are shown in table 3. The depth for the system - HOOP is $12$, and the depth for HOOP-MCTS_1 is $1$. From the table, we can see that the greedy approach of just looking ahead by 1 step is not enough to solve the rearrangement problem. This is because, when we look ahead only 1 step, we do not get enough reward/feedback from the world about what is a good path to take and hence makes it extremely hard to solve the problem. We can also see that, in the problems it does solve, it takes significantly longer to solve.
Table 3: Performance metrics for our method across different depths for MCTS search. Metrics: Success Score ($S$), Object Success Rate (OSR), Task Actions (TA), execution Time in minutes, and number of objects initially visible (#V). The difficulty parameters include the number of blocked goals (#BG), objects to be swapped (#Sw), blocking objects (#BP), and number of rooms (#Rm). NA: Not Applicable as no scenes were fully rearranged.
Dataset Objs #BG #Sw #BP #Rm #V HOOP HOOP-MCSTS_1
@@
Multi RoomR 10 1 1 0 2 2-3 32 65 710 7.89 0 7 NA
2 1 1 2 2-3 21 49 789 8.98 0 3 NA
10 2 1 0 3-4 1-2 30 62 1189 13.45 0 5 NA
2 1 1 3-4 1-2 18 44 1321 15.97 0 2 NA
15 1 1 0 3-4 2-3 22 59 1228 19.89 0 0 NA
2 1 1 3-4 2-3 14 41 1416 22.12 0 0 NA
20 2 1 0 3-4 2-4 17 55 1621 27.61 0 0 NA
2 1 1 3-4 2-4 10 36 1786 29.79 0 0 NA
Table 4: Comparison on RoomR dataset
Method HOOP Mirakhor et al. [2024a] Mirakhor et al. [2024b]
[Table 2] [Table 4]
Scene Success 49 43 34
We have run our system on the RoomR dataset, and we have results from Mirakhor et al. [2024a] and Mirakhor et al. [2024b] on the RoomR dataset Batra et al. [2020] and show the results in table 4 We can see that our success rate is slightly higher but this is with the caveat that we use more information at the beginning of rearrangement about what objects need to be moved.
A.5 Low-level RL policy
A.5.1 RL training
We trained our model using PPO [Schulman et al., 2017] with learning rate $\alpha=2.5\times 10^{-4}$, clip parameter $\epsilon=0.1$, value loss coefficient $c_{1}=0.5$ entropy coefficient $c_{2}=0.01$, GAE, recurrent policy, linear learning rate decay, 128 steps per update, 4 mini-batches. We train for 5 million steps on RTX 3080 GPU for 2 days. Observation Space (environment returns these at each step): RGB and Depth image. Agent position ($M^{2D}$ location + pitch + yaw). Object pick location for Pick policy training and object place location for training place policy. We also get information if an action succeeded or failed from the AI2Thor environment. The training process for the pick model involves randomly positioning the target object within a specified proximity to the agent. The goal is to pick a selected object successfully. For the place model, the training methodology follows a similar approach, with the key distinction being the absence of object detection requirements, as the agent begins each scenario already holding the object. In all training instances for the place model, the initial state consists of the agent holding an object, and the task involves depositing the object at a predetermined location.
A.5.2 Action space:
MoveAhead, MoveBack, MoveLeft, MoveRight LookUp,LookDown, RotateLeft,RotateRight, PickObject (for Pick policy only), PlaceObject (for Place policy only)
We allow the Pick and Place policies to have navigation actions because we might not always be in the perfect position to interact with an object and we want our policy to be able to handle these scenarios.
A.5.3 Reward Function:
Reward function: -1 for each navigation action, +50 for successful interaction action, -50 for failed interaction.
A.5.4 Transition function:
The transition function in AI2-Thor is deterministic. For the navigation actions, a move action moves the agent in the selected direction by 1 unit. Rotation action rotates the agent by 45 degrees in the given direction. Look action titles the head of the agent by 30 degrees in the given direction.
A.6 Frontier Exploration + hand coded interaction Method Details
A.6.1 Overview of the Heuristic Approach
Here, we provide a deeper dive into the construction of the hand coded heuristic method. The heuristic method serves as an important baseline that replaces the sophisticated MCTS search planning of our HOO-POMDP with hand-coded expert strategies while retaining the same belief update apparatus. This approach helps us evaluate the specific contribution of principled planning in handling uncertainty during rearrangement tasks.
Input: Environment $E$, Set of target objects $O$, Confidence threshold $\theta$ (default: 0.7)
Output: Rearranged objects at goal locations
Initialize belief state $B$ over object locations;
Initialize 2D occupancy grid map $M$ (initially all cells unknown);
Initialize empty set of frontier clusters $F$;
Initialize empty set of found objects Found;
while not all objects in $O$ are rearranged do
   Update belief state $B$ based on current observations;
 $F\leftarrow\text{IdentifyFrontierClusters}(M)$;
 foreach object $o\in O$ not yet rearranged do
    if $\max(B(o))>\theta$ and $o\notin\text{Found}$ then
         Add $o$ to Found;
 if $\text{Found}\neq\emptyset$ then
    $o_{\text{closest}}\leftarrow\text{GetClosestObjectFrom}(\text{Found})$;
      Remove $o_{\text{closest}}$ from Found;
    $\text{loc}_{\text{object}}\leftarrow\arg\max(B(o_{\text{closest}}))$;
    $\text{NavigateTo}(\text{loc}_{\text{object}})$;
    $\text{PickPlace}(o_{closest})$
 else
    $f_{\text{closest}}\leftarrow\text{GetClosestFrontierCluster}(F)$;
    $\text{NavigateTo}(f_{\text{closest}})$;
      Update $M$ based on new observations;
Algorithm 4 Hand-Coded Heuristic Method for Object Rearrangement
The algorithm presents our hand-coded heuristic method for object rearrangement tasks, serving as a critical baseline for comparison with the HOO-POMDP approach. As shown in Algorithm 4, this method maintains the same belief update apparatus while substituting sophisticated MCTS planning with expert-designed strategy.
The agent alternates between two phases: exploration (lines 21-23) and object interaction (lines 12-19). During exploration, the agent identifies frontier clusters at the boundary between known and unknown space, navigating to the closest frontier to systematically discover objects. The object interaction phase is triggered when the belief probability for any object exceeds the confidence threshold $\theta$ (line 9, 70% for our experiments). When this occurs, the agent temporarily halts exploration to retrieve and place the object at its goal location. The navigation system uses A* and the PickPlace policy is the same as in our HOOP system (uses RL pick, A* navigation and RL place policy to achieve PickPlace).
This approach provides a direct evaluation of the contribution of principled planning in handling uncertainty during rearrangement tasks, as it isolates the planning component while maintaining identical belief representations.
Since this is a simple system,it is limited to basic rearrangement and cannot be generalized to more complex problems - such as blocked goals. We would have to hand-design a new policy for each new case whereas our HOOP system can handle and plan for new scenarios much more easily.
A.6.2 Confidence Threshold Selection
The 70% confidence threshold was carefully calibrated for the specific characteristics of our problem:
  1. 1.
    Higher thresholds (>>70%): With our imperfect object detector (50-60% success rate), setting a higher confidence threshold would result in many objects never exceeding the threshold even when directly observed multiple times. This would lead to excessive exploration and low task completion rates.
  2. 2.
    Lower thresholds (<<70%): Setting a lower threshold increases the risk of false positives, where the agent attempts to interact with an object that isn’t actually present at the believed location. Our implementation treats failed pick attempts as definitive evidence that the object is not present, and the agent will not try again at that location. Therefore, false positives can permanently prevent successful rearrangement of certain objects.
  3. 3.
    Empirical optimization: The 70% threshold represents an empirically determined balance that maximizes overall task completion while minimizing both excessive exploration and false positive interactions.
A.7 Component Glossary for HOO-POMDP Framework
This glossary provides standardized terminology for the key components of our Hierarchical Object-Oriented POMDP (HOO-POMDP) framework to help readers maintain a clear understanding throughout the paper.
A.7.1 Core System Components
Term Definition Algorithm 1 Reference
HOO-POMDP Framework The complete hierarchical planning system for object rearrangement in partially observable environments. Full Algorithm 1
Perception Subsystem Processes RGB and depth images to detect objects and generate observations. PERCEPTIONSYSTEM() (line 8)
Belief Update Subsystem Maintains and updates probability distributions over possible object locations. BELIEFUPDATE() (line 9)
Fully described in Algorithm 2
Abstraction System Converts continuous low-level belief state into discrete abstract state. GENERATEABSSTATE() (line 10)
Abstract OOPOMDP Planner Uses POUCT to search through abstract actions and find the best sub-goal. POUCTPLANNER() (line 11)
Policy Executor Converts abstract sub-goals into sequence of executable low-level actions. GETLOWLEVELPOLICY() (line 14)
lowLevelPolicy.GetAct (Line 15)
It also executes the first low-level action in the above sequence in the environment. ENV.EXECUTE() (line 7)
A.7.2 Key Data Structures
Term Definition Representation
Belief State Probability distribution over possible locations for each object. $b=\prod_{i=1}^{n}b_{i}$
Abstract State Discrete representation of the world state used by the planner. $s=(s_{r},s_{targets})$
Observation Detection of object locations from the perception system. $z=(z_{robot},z_{objects})$
2D Occupancy Map Discretized grid representation of the navigable space. $M^{2D}$
A.7.3 Key Actions and Policies
Term Definition Examples
Abstract Actions High-level actions output by the abstract planner. MoveAB, Rotateangle
Sub-goals Task-oriented goal states for low-level policies to achieve. PickPlaceObjecti-goalloc
Low-level Actions Primitive actions executed directly in the environment. MoveAhead, PickObject
Low-level Policies Controllers that translate abstract actions into action sequences. Move, Rotate, PickPlace
A.7.4 Algorithm Components
Term Definition Reference
POUCT Partially Observable UCT - search algorithm for the abstract planner. Algorithm 3
Frontier Exploration Strategy to systematically explore unknown areas. Appendix A.6
A Algorithm Path planning algorithm for navigation between locations. Policy Executor (line 248)
A.7.5 Object State Representation
Term Definition Components
Object State Complete representation of an object in the abstract state. $(loc_{i},pick_{i},placeloc_{s},$
$is\_held,at\_goal,g_{i})$
Pick Location Location from which an object can be picked up. $pick_{i}$
Place Locations Set of locations where an object can be placed. $placeloc_{s}$
Goal Location Target location for task completion. $g_{i}$